we were asked to find the time taken for a ball to reach a specific point B. The abscissa of point B is 2m. the acceleration is -4, the initial velocity is 5, and the velocity at B is 3. I used the formula from the time equations that says that x=at^2/2+V initial * t + x initial. i got two answers for time, one which is 2 s and the other is 0.5 s. My teacher used another way and got only 0.5 seconds. Why is 2 seconds rejected?
1
Expert's answer
2019-11-28T08:29:19-0500

found time values must satisfy the equation for speed

"v=v_0+a \cdot t"

By condition

"3=5-4 \cdot t"

equality will be fulfilled only when "t_2=0.5s"

"3=5-4 \cdot t_2"

"3=5-4 \cdot 0.5"

"3=3"